# Rethinking Multi-domain Generalization with A General Learning Objective

### v1

@@ Abstract
Multi-domain generalization (mDG) is universally aimed to minimize the discrepancy between training and testing distributions to enhance marginal-to-label distribution mapping. However, existing mDG literature lacks a general learning objective paradigm and often imposes constraints on static target marginal distributions. In this paper, we propose to leverage a ${\mathbf{Y}}$-mapping to relax the constraint. We rethink the learning objective for mDG and design a new general learning objective to interpret and analyze most existing mDG wisdom. This general objective is bifurcated into two synergistic amis: learning domain-independent conditional features and maximizing a posterior. Explorations also extend to two effective regularization terms that incorporate prior information and suppress invalid causality, alleviating the issues that come with relaxed constraints. We theoretically contribute an upper bound for the domain alignment of domain-independent conditional features, disclosing that many previous mDG endeavors actually optimize partially the objective and thus lead to limited performance. As such, our study distills a general learning objective into four practical components, providing a general, robust, and flexible mechanism to handle complex domain shifts. Extensive empirical results indicate that the proposed objective with ${\mathbf{Y}}$-mapping leads to substantially better mDG performance in various downstream tasks, including regression, segmentation, and classification.
@@ 1 Introduction
$d_{n}\in\mathcal{D},n\leq N$ ; $d^{\prime}\in\mathcal{D}$:
The $n$-th observed domains in all domains; Unseen domains in all domains.
${\mathbf{X}},{\mathbf{Y}}$; ${\mathbf{X}}_{n},{\mathbf{Y}}_{n}$; ${\mathbf{X}}^{\prime},{\mathbf{Y}}^{\prime}$.
All observations and targets; Observations and targets in $d_{n}$; Observations and targets in $d^{\prime}$.
$P(x)$:
Distributions where $x$ corresponds to the random variables.
$\phi,\psi$:
Learnable transformations that codify ${\mathbf{X}},{\mathbf{Y}}$ into the same latent RKHS.
$\phi({\mathbf{X}}),\psi({\mathbf{Y}})$:
Mapped ${\mathbf{X}},{\mathbf{Y}}$. Within the RKHS realm, $\phi({\mathbf{X}}),\psi({\mathbf{Y}})$ follow Multivariate Gaussian Distributions.
$\mathcal{O}$; $R(\cdot)$; $\sigma_{\cdot,\cdot}$:
Prior knowledge (oracle model); Empirical risks; Covariance between two variables.
$\mathcal{C}:\phi({\mathbf{X}}),\psi({\mathbf{Y}})\to{\mathbf{Y}}$:
Predictor that predicts ${\mathbf{Y}}$ from $\phi({\mathbf{X}}),\psi({\mathbf{Y}})$.
$D_{\mathrm{KL}}(\cdot\|\cdot);H(\cdot);H_{c}(\cdot,\cdot)$:
KL divergence; Entropy; Cross-entropy.
@@ 3.1 Theoretical Details
where $a\triangleq\sum\nolimits_{n=1}^{N}H(P({\phi({\mathbf{X}}_{n}),\psi({\mathbf{Y}}_{n})}))$ is constant w.r.t $\phi,\psi$, hence ignored during optimization. The novel PUB is derived from Eq. 6, is:
@@
where is GReg2. See more mathematical details in Supplementary 7.
Our experiments also unveil the phenomenon of invalid causality within invariant feature learning, where suppressing it could improve generalizability. The investigation of previous objectives also discloses that, in addressing the varying $P({\mathbf{Y}}|\mathcal{D})$, constructs akin to GReg2 are often implicitly included (see Table 1), though their efficacy was not explicitly stated. Furthermore, their efficacy may be compromised due to the lack of $\psi$ and other objective terms.
Then, we assume that $\phi(X),\psi(Y)$ in the RKSH follow Multivariate Gaussian-like Distributions which are denoted as $\mathcal{N}(\phi({\mathbf{X}});\mu_{\mathbf{X}},\Sigma_{{\mathbf{X}}{\mathbf{X}}}),\mathcal{N}(\psi({\mathbf{Y}});\mu_{\mathbf{Y}},\Sigma_{{\mathbf{Y}}{\mathbf{Y}}})$. Then, $P(\phi({\mathbf{X}})|\psi({\mathbf{Y}}))$ follows $\mathcal{N}(\phi({\mathbf{X}})|\psi({\mathbf{Y}});\mu_{{\mathbf{X}}|{\mathbf{Y}}},\Sigma_{{\mathbf{X}}{\mathbf{X}}|{\mathbf{Y}}})$. GReg2 can be simplified as:
@@
where the inequality stands owing to the Condition Reducing Entropy. This implies $H(\mathcal{N}(\phi({\mathbf{X}});\mu_{\mathbf{X}},\Sigma_{{\mathbf{X}}{\mathbf{X}}}))\geq H(\mathcal{N}(\phi({\mathbf{X}})\mid\psi({\mathbf{Y}});\mu_{{\mathbf{X}}|{\mathbf{Y}}},\Sigma_{{\mathbf{X}}{\mathbf{X}}|{\mathbf{Y}}}))$, deduced from $|\Sigma_{{\mathbf{X}}{\mathbf{X}}}|\geq|\Sigma_{{\mathbf{X}}{\mathbf{X}}|{\mathbf{Y}}}|\geq 0$, considering they are positive semi-definite. Thus, minimization of Eq. 10 occurs iff $|\Sigma_{{\mathbf{X}}{\mathbf{X}}}|=|\Sigma_{{\mathbf{X}}{\mathbf{X}}|{\mathbf{Y}}}|$, reformulating the task as $min_{\phi,\psi}|\Sigma_{{\mathbf{X}}{\mathbf{X}}}|-|\Sigma_{{\mathbf{X}}{\mathbf{X}}|{\mathbf{Y}}}|$, where $\Sigma_{{\mathbf{X}}{\mathbf{X}}\mid{\mathbf{Y}}}=\Sigma_{{\mathbf{X}}{\mathbf{X}}}-\Sigma_{{\mathbf{X}}{\mathbf{Y}}}\Sigma_{{\mathbf{Y}}{\mathbf{Y}}}^{-1}\Sigma_{{\mathbf{Y}}{\mathbf{X}}}$, per [20]. Therefore, GReg2 is simplified as minimizing Conditional Feature Shift (CFS):
@@
3.2 Empirical Losses Derivation.
@@
To mitigate domain shifts and learn domain invariance, we minimize cross-domain conditional feature distribution discrepancies. Specifically, the mean and variance of the joint distribution of $(\phi({\mathbf{X}}),\psi({\mathbf{Y}}))$ in each domain are estimated using VAE encoders. Consider $n$-pairs means and variance of $n$ domains, we derive a joint Gaussian distribution expression $P(\phi({\mathbf{X}}_{n}),\psi({\mathbf{Y}}_{n}))\triangleq\mathcal{N}({\bm{x}}_{n},{\bm{y}}_{n};\mu_{n},\Sigma_{n})$. Accordingly, we establish $\mathbb{E}[P({\phi({\mathbf{X}}_{n}),\psi({\mathbf{Y}}_{n})})]\triangleq\mathcal{N}(\bar{{\bm{x}}},\bar{{\bm{y}}};\bar{\mu},\bar{\Sigma})$ where $\bar{\mu}=\mathbb{E}[\mu_{n}],\bar{\Sigma}=\mathbb{E}[\Sigma_{n}]$. Base on PUB in Eq. 7, we introduce $\mathcal{L}_{A1}$ to minimize the conditional feature gap across domains:
@@
where $\Sigma_{{\mathbf{X}}{\mathbf{Y}}}=\mathbb{E}[(\phi({\mathbf{X}})-\mathbb{E}[\phi({\mathbf{X}})])^{\top}(\phi({\mathbf{Y}})-\mathbb{E}[\phi({\mathbf{Y}})])]$, and a similar calculation process is done for $\Sigma_{{\mathbf{Y}}{\mathbf{Y}}}$ and $\Sigma_{{\mathbf{Y}}{\mathbf{X}}}$. The final loss is a weighted combination of the above losses:
@@
Detailed hyper-parameters settings can be seen in Supplementary 10.
@@ 4 Connection to Previous Methods
Effect of oracle model $\mathcal{O}$. As stated by MIRO [18] and SIMPLE [29], a generalized $\mathcal{O}$ comprising both seen and unseen domains yields significant improvements. During the derivation of Eq. 7, we find that the disregard GAim1 term in MIRO [18] and SIMPLE [29] may result in inferior outcomes to our proposed objective.
@@
For readability, we have divided the overall mDG objective into four aspects despite all terms being interconnected. Specifically, as shown by PUB in Eq. 7, GReg1 collaborating with GAim1 brings more performance gains than the case when it is solely applied. Moreover, Eq. 8 shows that the side effect of invalid causality in GAim1 is alleviated by combining with GReg2, underscoring the significance of combining learning invariance, integrating prior knowledge, and suppressing invalid causally. It also suggests that all terms are synergistic and contribute together to improved results.
@@ 5.1 Toy experiments on synthetic datasets
Results. Toy experiment results are reported in Table 3, which are also visualized in Figure 3. It is observed that across all settings, employing $\psi$ with $\mathcal{L}_{A1}$ yields superior results, outperforming ERM and ERM+$\mathcal{L}_{A1}$($\phi$) without $\psi$, validating the enhanced generalization effect brought by utilizing $\psi$ whenever $Y$ varies per domain, supporting Eq. 17. Figure 3 shows that $\psi$ does learn the invariant representations for $Y$ to relax previous $Y$-invariant assumption. Specifically, learning the invariance of $Y$ with $\psi$ results in superior invariant representations as the latent representations of $X,Y$ are primarily linear, aligning with $X$ and $Y$’s linear relationship during data construction. The bottom-left figures reveal that though ERM has learned the most invariant $\phi(X)$, it suffers the worst test loss, indicating that a well-learned invariant $\phi(X)$ is not sufficient when $Y$ also has domain-dependent traits. The results also suggest that assuming that $Y$ vary across domains, using $\mathcal{L}_{A1}$ without $\psi$ may not yield superior results.
@@ 5.3 Segmentation on benchmark datasets
Results. Table 6 shows the efficacy of our proposed objective in segmentation tasks upon introducing $\psi$. Ablation results highlight that using $\psi$ alongside GAim1 can enhance baseline performance, experimentally substantiating that the introduction of $\psi$, in relaxing assumptions, boosts performance for better generalization. Using GReg1 alone also improves average mIoU. Importantly, the most enhancement in average mIoU is observed when GReg1 and GAim1 are used together, which finds validation in the PUB derivation in Eq. 7. See more results and visualizations in Supplementary 11.
@@ 5.5 Ablation studies
To better compare our objective with previous objectives, we conduct a systematic ablation study on the classification task since most previous objectives are only available for classification due to the lack of $\psi$. Experimental setup. In the ablation studies, we test varied terms (see Table 8) combinations on the HomeOffice dataset using SWAG pre-training [46] and SWAD [6]. Every experiment is repeated in three trials, sharing the same hyper-parameter settings for evaluation. See Supplementary 10 for more experimental details.
@@
GAim2 (ERM)
78.4±0.7 68.3±0.5 85.8±0.4 85.8±0.3 79.6±0.2 0.0
GAim2 + iAim1 (DANN)
79.1±1.0 68.6±0.0 85.6±0.8 86.1±0.5 79.8±0.2 +0.2
GAim2 + GAim1 (CDANN, CIDG)
79.1±0.7 69.1±0.1 85.7±0.5 86.3±0.6 79.9±0.4 +0.3
GAim2 +iReg2 (CORAL+$\psi$)
79.1±0.1 69.9±0.4 86.0±0.1 86.3±0.4 80.3±0.2 +0.7
@@
GAim2 + GAim1
+ GReg2 (MDA+$\psi$)
79.5±1.1 69.2±1.2 86.2±0.2 86.5±0.2 80.3±0.0 +0.7
@@
GAim2 + GReg1 (MIRO, SIMPLE)
83.2±0.6 72.6±1.1 89.9±0.5 90.2±0.1 84.0±0.2 0.0
@@
GAim2 + GReg1
+ GAim1 + GReg2 (GMDG)
84.1±0.2 74.3±0.9 89.9±0.4 90.6±0.1 84.7±0.2 +0.7
@@
3). Learning invariance is crucial, regardless of whether integrating prior knowledge. Evidently, learning invariance facilitates improvement whether prior is applied or not, as validated in the PUB derivation in Eq. 7. This contradicts MIRO’s argument that achieving similar representations to a prior can replace the need for learning invariance.

### v2

@@ Abstract
Multi-domain generalization (mDG) is universally aimed to minimize the discrepancy between training and testing distributions to enhance marginal-to-label distribution mapping. However, existing mDG literature lacks a general learning objective paradigm and often imposes constraints on static target marginal distributions. In this paper, we propose to leverage a ${\mathbf{Y}}$-mapping to relax the constraint. We rethink the learning objective for mDG and design a new general learning objective to interpret and analyze most existing mDG wisdom. This general objective is bifurcated into two synergistic amis: learning domain-independent conditional features and maximizing a posterior. Explorations also extend to two effective regularization terms that incorporate prior information and suppress invalid causality, alleviating the issues that come with relaxed constraints. We theoretically contribute an upper bound for the domain alignment of domain-independent conditional features, disclosing that many previous mDG endeavors actually optimize partially the objective and thus lead to limited performance. As such, our study distills a general learning objective into four practical components, providing a general, robust, and flexible mechanism to handle complex domain shifts. Extensive empirical results indicate that the proposed objective with ${\mathbf{Y}}$-mapping leads to substantially better mDG performance in various downstream tasks, including regression, segmentation, and classification. Code is available at https://github.com/zhaorui-tan/GMDG/tree/main.
@@ 1 Introduction
$d_{n}\in\mathcal{D},n\leq N$ ; $d^{\prime}\in\mathcal{D}$: The $n$-th observed domains in all domains; Unseen domains in all domains.
${\mathbf{X}},{\mathbf{Y}}$; ${\mathbf{X}}_{n},{\mathbf{Y}}_{n}$; ${\mathbf{X}}^{\prime},{\mathbf{Y}}^{\prime}$. All observations and targets; Observations and targets in $d_{n}$; Observations and targets in $d^{\prime}$.
$P(x)$: Distributions where $x$ corresponds to the random variables.
$\phi,\psi$: Learnable transformations that codify ${\mathbf{X}},{\mathbf{Y}}$ into the same latent RKHS.
$\phi({\mathbf{X}}),\psi({\mathbf{Y}})$: Mapped ${\mathbf{X}},{\mathbf{Y}}$. Within the RKHS realm, $\phi({\mathbf{X}}),\psi({\mathbf{Y}})$ follow Multivariate Gaussian Distributions.
$\mathcal{O}$; $R(\cdot)$; $\sigma_{\cdot,\cdot}$: Prior knowledge (oracle model); Empirical risks; Covariance between two variables.
$\mathcal{C}:\phi({\mathbf{X}}),\psi({\mathbf{Y}})\to{\mathbf{Y}}$: Predictor that predicts ${\mathbf{Y}}$ from $\phi({\mathbf{X}}),\psi({\mathbf{Y}})$.
$D_{\mathrm{KL}}(\cdot\|\cdot);H(\cdot);H_{c}(\cdot,\cdot)$: KL divergence; Entropy; Cross-entropy.
@@ 3.1 Theoretical Details
where $a\triangleq\sum\nolimits_{n=1}^{N}H(P({\phi({\mathbf{X}}_{n}),\psi({\mathbf{Y}}_{n})}))$ is constant w.r.t $\phi,\psi$, hence ignored during optimization. The novel PUB is derived from Eq. LABEL:eq:bound2, is:
@@
where is GReg2. See more mathematical details in Supplementary 7. Our experiments also unveil the phenomenon of invalid causality within invariant feature learning, where suppressing it could improve generalizability. The investigation of previous objectives also discloses that, in addressing the varying $P({\mathbf{Y}}|\mathcal{D})$, constructs akin to GReg2 are often implicitly included (see Table 1), though their efficacy was not explicitly stated. Moreover, their efficacy may be compromised due to the lack of $\psi$ and other objective terms.
Then, we assume that $\phi(X),\psi(Y)$ in the RKSH follow Multivariate Gaussian-like Distributions which are denoted as $\mathcal{N}(\phi({\mathbf{X}});\mu_{\mathbf{X}},\Sigma_{{\mathbf{X}}{\mathbf{X}}}),\mathcal{N}(\psi({\mathbf{Y}});\mu_{\mathbf{Y}},\Sigma_{{\mathbf{Y}}{\mathbf{Y}}})$. $P(\phi({\mathbf{X}})|\psi({\mathbf{Y}}))$ follows $\mathcal{N}(\phi({\mathbf{X}})|\psi({\mathbf{Y}});\mu_{{\mathbf{X}}|{\mathbf{Y}}},\Sigma_{{\mathbf{X}}{\mathbf{X}}|{\mathbf{Y}}})$. GReg2 can be simplified as:
@@
where the inequality stands owing to the Condition Reducing Entropy. This implies $H(\mathcal{N}(\phi({\mathbf{X}});\mu_{\mathbf{X}},\Sigma_{{\mathbf{X}}{\mathbf{X}}}))\!\!\geq\!\!H(\mathcal{N}(\phi({\mathbf{X}})\mid\psi({\mathbf{Y}});\mu_{{\mathbf{X}}|{\mathbf{Y}}},\Sigma_{{\mathbf{X}}{\mathbf{X}}|{\mathbf{Y}}}))$, deduced from $|\Sigma_{{\mathbf{X}}{\mathbf{X}}}|\!\!\geq\!\!|\Sigma_{{\mathbf{X}}{\mathbf{X}}|{\mathbf{Y}}}|\!\!\geq\!\!0$, considering they are positive semi-definite. Distinct from the [52, 18] which decompose causal effects through extra networks, our method is based on transfer entropy (TE) by ensuring $TE(\phi({\mathbf{X}})\!\!\!\to\!\!\!\psi(Y))\!\!\!\geq\!\!\!TE(\phi({\mathbf{X}})\!\!\!\to\!\!\!\psi(Y))$, i.e., $H(\phi({\mathbf{X}})|\psi(Y))\!\!\!\geq\!\!\!H(\psi(Y)|\phi({\mathbf{X}})).$ Thus, minimization of Eq. LABEL:eq:cfs_detial occurs iff $|\Sigma_{{\mathbf{X}}{\mathbf{X}}}|=|\Sigma_{{\mathbf{X}}{\mathbf{X}}|{\mathbf{Y}}}|$, reformulating the task as $min_{\phi,\psi}|\Sigma_{{\mathbf{X}}{\mathbf{X}}}|\!\!\!-\!\!\!|\Sigma_{{\mathbf{X}}{\mathbf{X}}|{\mathbf{Y}}}|$, where $\Sigma_{{\mathbf{X}}{\mathbf{X}}\mid{\mathbf{Y}}}=\Sigma_{{\mathbf{X}}{\mathbf{X}}}\!\!-\!\!\Sigma_{{\mathbf{X}}{\mathbf{Y}}}\Sigma_{{\mathbf{Y}}{\mathbf{Y}}}^{-1}\Sigma_{{\mathbf{Y}}{\mathbf{X}}}$, per [21]. Therefore, GReg2 is simplified as minimizing Conditional Feature Shift (CFS):
@@
3.2 Empirical losses derivations
@@
To mitigate domain shifts and learn domain invariance, we minimize cross-domain conditional feature distribution discrepancies. Specifically, the mean and variance of the joint distribution of $(\phi({\mathbf{X}}),\psi({\mathbf{Y}}))$ in each domain are estimated using VAE encoders. Consider $n$-pairs means and variance of $n$ domains, we derive a joint Gaussian distribution expression $P(\phi({\mathbf{X}}_{n}),\psi({\mathbf{Y}}_{n}))\triangleq\mathcal{N}({\bm{x}}_{n},{\bm{y}}_{n};\mu_{n},\Sigma_{n})$. Accordingly, we establish $\mathbb{E}[P({\phi({\mathbf{X}}_{n}),\psi({\mathbf{Y}}_{n})})]\triangleq\mathcal{N}(\bar{{\bm{x}}},\bar{{\bm{y}}};\bar{\mu},\bar{\Sigma})$ where $\bar{\mu}=\mathbb{E}[\mu_{n}],\bar{\Sigma}=\mathbb{E}[\Sigma_{n}]$. Base on PUB in Eq. LABEL:eq:PUB, we introduce $\mathcal{L}_{A1}$ to minimize the conditional feature gap across domains:
@@
where $\Sigma_{{\mathbf{X}}{\mathbf{Y}}}=\mathbb{E}[(\phi({\mathbf{X}})-\mathbb{E}[\phi({\mathbf{X}})])^{\top}(\phi({\mathbf{Y}})-\mathbb{E}[\phi({\mathbf{Y}})])]$, and a similar calculation process is done for $\Sigma_{{\mathbf{Y}}{\mathbf{Y}}}$ and $\Sigma_{{\mathbf{Y}}{\mathbf{X}}}$. The final loss is a weighted combination of the above losses [^1] See detailed hyper-parameters settings in Supplementary 10.:
@@ 4 Connection to previous methods
Effect of oracle model $\mathcal{O}$. As stated by MIRO [19] and SIMPLE [30], a generalized $\mathcal{O}$ comprising both seen and unseen domains yields significant improvements. During the derivation of Eq. LABEL:eq:PUB, we find that the disregard GAim1 term in MIRO [19] and SIMPLE [30] may result in inferior outcomes to our proposed objective.
@@
For readability, we have divided the overall mDG objective into four aspects despite all terms being interconnected. Specifically, as shown by PUB in Eq. LABEL:eq:PUB, GReg1 collaborating with GAim1 brings more performance gains than the case when it is solely applied. Moreover, Eq. 8 shows that the side effect of invalid causality in GAim1 is alleviated by combining with GReg2, underscoring the significance of combining learning invariance, integrating prior knowledge, and suppressing invalid causally. It also suggests that all terms are synergistic and contribute together to improved results.
@@ 5.1 Toy experiments on synthetic datasets
Results. Toy experiment results are reported in Table 3, which are also visualized in Figure 4. It is observed that across all settings, employing $\psi$ with $\mathcal{L}_{A1}$ yields superior results, outperforming ERM and ERM+$\mathcal{L}_{A1}$($\phi$) without $\psi$, validating the enhanced generalization effect brought by utilizing $\psi$ whenever $Y$ varies per domain, supporting Eq. 17. Supplementary Figure 4 shows that $\psi$ does learn the invariant representations for $Y$ to relax previous $Y$-invariant assumption. Specifically, learning the invariance of $Y$ with $\psi$ results in superior invariant representations as the latent representations of $X,Y$ are primarily linear, aligning with $X$ and $Y$’s linear relationship during data construction. The bottom-left figures reveal that though ERM has learned the most invariant $\phi(X)$, it suffers the worst test loss, indicating that a well-learned invariant $\phi(X)$ is not sufficient when $Y$ also has domain-dependent traits. The results also suggest that assuming that $Y$ vary across domains, using $\mathcal{L}_{A1}$ without $\psi$ may not yield superior results.
@@ 5.3 Segmentation on benchmark datasets
Results. Table 6 shows the efficacy of our proposed objective in segmentation tasks upon introducing $\psi$. Ablation results highlight that using $\psi$ alongside GAim1 can enhance baseline performance, experimentally substantiating that the introduction of $\psi$, in relaxing assumptions, boosts performance for better generalization. Using GReg1 alone also improves average mIoU. Importantly, the most enhancement in average mIoU is observed when GReg1 and GAim1 are used together, which finds validation in the PUB derivation in Eq. LABEL:eq:PUB. See more results and visualizations in Supplementary 11.
@@ 5.5 Ablation studies
To better compare our objective with previous objectives, we conduct a systematic ablation study on the classification task since most previous objectives are only available for classification due to the lack of $\psi$. Experimental setup. In the ablation studies, we test varied terms (see Table 8) combinations on the HomeOffice dataset using SWAG pre-training [47] and SWAD [6]. Every experiment is repeated in three trials, sharing the same hyper-parameter settings for evaluation. See Supplementary 10 for more details.
@@
GAim2 (ERM) 78.4±0.7 68.3±0.5 85.8±0.4 85.8±0.3 79.6±0.2 0.0
GAim2 + iAim1 (DANN) 79.1±1.0 68.6±0.0 85.6±0.8 86.1±0.5 79.8±0.2 +0.2
GAim2 + GAim1 (CDANN, CIDG) 79.1±0.7 69.1±0.1 85.7±0.5 86.3±0.6 79.9±0.4 +0.3
GAim2 +iReg2 (CORAL+$\psi$) 79.1±0.1 69.9±0.4 86.0±0.1 86.3±0.4 80.3±0.2 +0.7
@@
GAim2 + GAim1 + GReg2 (MDA+$\psi$) 79.5±1.1 69.2±1.2 86.2±0.2 86.5±0.2 80.3±0.0 +0.7
@@
GAim2 + GReg1 (MIRO, SIMPLE) 83.2±0.6 72.6±1.1 89.9±0.5 90.2±0.1 84.0±0.2 0.0
@@
GAim2 + GReg1 + GAim1 + GReg2 (GMDG) 84.1±0.2 74.3±0.9 89.9±0.4 90.6±0.1 84.7±0.2 +0.7
@@
3). Learning invariance is crucial, regardless of whether integrating prior knowledge. Evidently, learning invariance facilitates improvement whether prior is applied or not, as validated in the PUB derivation in Eq. LABEL:eq:PUB. This contradicts MIRO’s argument that achieving similar representations to a prior can replace the need for learning invariance.